An electric eel develops a potential difference of 470 V , driving a current of 0.85 A for a 1.0 ms pulse.

1. Find the power of this pulse.

2. Find the total energy of this pulse.

3. Find the total charge that flows during the pulse.
Answer #1

1) Power = V*I

= 470*0.85

= 399.5 W

2) Total energy of the pulse = Power*time

= 399.5*1*10^-3

= 0.3995 J

3) I = dQ/dt

==> dQ = I*dt

= 0.85*(1*10^-3)

= 8.5*10^-4 C